17 September 2010 2nd half

  • LBC Chip 2 still dead. See lbcr.20100917.071852.fits. It is not responding to starlight, but has only bias counts plus crosstalk from chip 1.
    • This is necessitating a change in dofpia. We are making it use chip 3 instead. No success yet in modifying the IDL code. Message from Vincenzo in the meantime.
  • 00:45: Vincenzo suggested checking video & bias cable connections. Must turn off everything before. Turning off cameras & getting telescope/dome ready.
  • 01:45: Went up to the Red cryostat and saw that the shield on the cable to chip 2 was broken. Checked connection. Will try again.
  • 02:00: 090213 R still shows chip 2 dead.
  • 02:30: Fernando suggested that the cable be removed & examined. Nothing we can do for it right now, but will remove it in the morning
and then depending on who is up tomorrow in the day, will either keep it here or bring it to town.
  • 02:40: chip 3 used to collimate R. Obtaining superfoc data rV.
  • 02:44: check focal plane tilt. Both cryostats were removed;
    • superfoc PA=0 0944 - 0950.
    • superfoc PA=180
    • superfoc PA=180
    • superfoc PA = 0
  • 03:15 check guiding. Red master/Blue slave. 101605/2 B/R (these have red master/blue slave)
    • check first image - rotation.
    • check tilt from these images also
    • guiding looks OK.
  • 03:30 Now switch to have Blue master/Red slave. 1033...
    • guiding looks OK by eye
    • meanwhile Fernando called to propose a new test.
  • 03:50 Going to test rotator positioning on first dither. In past rotator did not move all the way for the first dither.
    • testfirstdither.ob uses SDT_Uspec & i-SLOAN. We note that the first dither has PA = 199.97 and whlie it is prepping for the next, it moves rotator.
  • 04:16 Checking co-pointing with mirradjust IRAF script.
    • slew to near zenith: am=1.002, B/R 112328/112325
    • result (hardwired mirradjust to work on chip 1 --- note chip is a parameter, but the 'official' version is hardwired to chip 2.
The following values are _offsets_ which need to be applied to the 
SX Global Offsets in PSF:
dX = 3.00 mm, dY = -1.74 mm, dRX = -36.51 arcsec and dRY = -63.07 arcsec

Entered tip = -37 and tilt = -63. (=after discussion, change this to tilt = +63, not -63... tilt corresponds to -dRY while tip corresponds to +dRX).
  • 113954, 113952 B,R after tip/tilt changed.
    • gives 3.9, -6 as tip/tilt

  • slew to low elevation to check limits. We his M1 X limit was hit. Both mirrors were tilted together. We reduced tilt to +23 and put tilt = -43 onto DX (added later after discussion with JH)
  • back to zenith. Collimation held.
  • now back to low elevation: 27 deg. 115825,115821 B,R

  • 05:33: doing Fernando's experiment: cable 3 is plugged into chip 2. lbcr.20100917.123346.fits. Chip 3 cable plugged into chip 2. Now what display tool calls chip 3 is alive.
  • Took a longer exposure also - 123803

  • To reproduce co-pointing for next nights, at the start of the night, we must give tip = -37 and tilt = +23 to SX. (to DX we should give tilt = -43).

-- OlgaKuhn - 17 Sep 2010
Topic revision: r2 - 17 Sep 2010, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback